Quantitative Problem 1: Suppose that 1 Swedish krona could be purchased in the foreign exchange market today for $0.12. If the krona appreciated 10% tomorrow against the dollar, how many kronas would a dollar buy tomorrow? Do not round intermediate calculations. Round your answer to two decimal places.

Quantitative Problem 2: Suppose the exchange rate between the U.S. dollar and the South African rand was 16 rand = $1 and the exchange rate between the U.S. dollar and the Israeli shekel was 1 shekel = $0.30. What was the exchange rate between the South African rand and the Israeli shekel? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Dollars required for one Krona tomorrow = 0.12*(1+10%) = $0.132

Hence. kronas in one dollar = 1/0.132 = 7.5757 kronas

i.e. 7.58 kronas

2.Rand/Shekel = (Rand/Dollar)*(Dollar/Shekel)

= 16*0.30

= Rand 4.80/Shekel